in 2005,5.5 out of every 50 coast guard personnel were women, IF THERE ARE 40,639 TOTAL COAST GUARD PERSONNEL ON ACTIVE DUTY, ESTIMATE THE NUMBER OF WOMEN. ROUND TO THE NEAREST WHOLE
Answer by Theo(13342)

5.5 out of every 50 coast guard personnel were women.

take 40,639 and divide it by 50 and you get 812.78 sets of 50.

since 5.5 out of each set of 50 are women, there must be 5.5 * 812.78 = 4470.29 women.

this rounds down to 4470 women.

since 5.5 out of each set of 50 are women, then 44.5 out of each set of 50 are men.

since 44.5 out of each set of 50 are men, there must be 44.5 * 812.78 = 36168.71 men.

this rounds up to 36169 men.

you have 4470 women and 36169 men for a total of 40639 personnel.

the ratio of women to total is 4470 / 40639 = .109992864.

multiply this by 50 and you get approximately 5.4996432 women out of every 50 personnel.

the difference between this and 5.5 out of every 50 is due to rounding.

likewise, the ratio of men to total is 36169 / 40639 = .890007136.

mjultiply this by 50 and you get approximately 44.5003568 men out of every 50 personnel.

the difference between this and 4.5 out of every 50 is, once again, due to rounding.